THIS IS THE SECOND TIME I HAVE REQUESTED THIS QUESTION. THE FIRST ANSWER IS THE ONLY ONE CORRECT Here are simplified financial statements for Phone Corporation in a recent year:INCOME STATEMENT(Figures in $ millions Net sales$12,600 Cost of goods sold3,700 Other expenses4,127 Depreciation2,338 Earnings before interest and taxes (EBIT)$2,375 Interest expense 655 Income before tax$1,720 Taxes (at 35%)602 Net income$1,118 Dividends$816 BALANCE SHEET(Figures in $ millions)End of YearStart ofYearAssets End of year start of the year Cash and marketable securities $83 $152 Receivables 2,082 2,370 Inventories 157 208 Other current assets 837 902 Total current assets 3,159 3,632 Net property, plant, and equipment 19,913 19,855 Other long-term assets 4156 3710 Total assets 27228 27197 Liabilities and shareholders’ equity Payables 2504 2980 Short-term deb 1389 1543 Other current liabilities 781 757 Total current liabilities 4674 5280 Long-term debt and leases 8512 7907 Other long-term liabilities 6118 6089 Shareholders’ equity 7924 7921 Total liabilities and shareholders’ equity $27,228 $27,197 Calculate the following financial ratios for Phone Corporation: (Use 365 days in a year. Do not round intermediate calculations. Round your final answers to 2 decimal places.) a.Return on equity (use average balance sheet figures) b.Return on assets (use average balance sheet figures) c.Return on capital (use average balance sheet figures) d.Days in inventory (use start-of-year balance sheet figures) e.Inventory turnover (use start-of-year balance sheet figures) f.Average collection period (use start-of-year balance sheet figures) g.Operating profit margin h.Long-term debt ratio (use end-of-year balance sheet figures) i.Total debt ratio (use end-of-year balance sheet figures) j.Times interest earned3 k.Cash coverage ratio l.Current ratio (use end-of-year balance sheet figures) m.Quick ratio (use end-of-year balance sheet figures)
Answer #1

As per rules I am answering the first 4 subparts of the question

A: ROE = Net Income /Average Equity

= 1118/ ((7924+7921)/1)

=14.11%

B: ROA = Net Income/ Average assets

= 1118/ ((27228+27197)/2)

=4.11%

C: Return on capital = EBIT/ (Total assets- Current liabilities)

= 2375/ (27228-4674)

=10.53%

D: Days in inventory = Inventory * 365/ COGS

= 208*365/ 3700

=20.52 days